# Non-Iterative Disentangled Unitary Coupled-Cluster based on Lie-algebraic structure

### v1

@@ Abstract
Due to their non-iterative nature, fixed unitary coupled cluster (UCC) ansätze are attractive for performing quantum chemistry variational quantum eigensolver (VQE) computations as they avoid pre-circuit measurements on a quantum computer. However, achieving chemical accuracy for strongly correlated systems with UCC requires further inclusion of higher-order fermionic excitations beyond triples increasing circuit depth. We introduce $k$-NI-DUCC, a fixed and non-iterative disentangled unitary coupled cluster compact ansatz, based on specific ``$k^{\prime\prime}$ sets of “qubit” excitations, eliminating the needs for fermionic-type excitations. These elements scale linearly ($\mathcal{O}(n)$) by leveraging Lie algebraic structures, with $n$ being the number of qubits. The key excitations are screened through specific selection criteria, including the enforcement of all symmetries, to ensure the construction of a robust set of generators. NI-DUCC employs ``$k^{\prime\prime}$ products of the exponential of $\mathcal{O}(n)$- anti-Hermitian Pauli operators, where each operator has a length $p$. This results in a fewer two-qubit CNOT gates circuit, $\mathcal{O}(knp)$, suitable for hardware implementations. Tested on LiH, H6 and BeH2, NI-DUCC-VQE achieves both chemical accuracy and rapid convergence even for molecules deviating significantly from equilibrium. It is hardware-efficient, reaching the exact full configuration interaction energy solution at specific layers, while reducing significantly the VQE optimization steps. While NI-DUCC-VQE effectively addresses the gradient measurement bottleneck of ADAPT-VQE-like iterative algorithms, the classical computational cost of constructing the $\mathcal{O}(n)$ set of excitations increases exponentially with the number of qubits. We provide a first implementation for constructing the generators’ set, able to handle up to 20 qubits and discuss the efficiency perspectives.
@@ Introduction
To address these issues, constructing an ansatz tailored to system-specific criteria offers a promising solution for achieving shallower circuits. The popular ADAPT-VQE iterative algorithm is a notable example of this strategy [32]. While capable of extracting electronic correlation energy using a small subset of operators, the operator selection process corresponds itself to a high computational overhead associated to gradients measurements. Despite subsequent efforts to minimize these costs, ADAPT-VQE versions [33, 34, 35, 36, 37, 38, 39, 40] continue to face gradient measurement challenges, especially as molecular size increases, while maintaining the necessary chemical accuracy. Similar to ADAPT-VQE methods, the Qubit coupled cluster (QCC) approach [41, 42] and its variants, including iterative Qubit coupled cluster (iQCC) [43], offer promising strategies for designing a compact wavefunction. In iQCC, generators are chosen from a specific set of operators, including all possible Pauli strings, to ensure energy reduction. These generators are sampled using first-order derivative calculations, resulting in a shallow circuit depth. Unlike ADAPT-VQE, the iQCC method avoids the need for full re-optimization of all parameters at each iteration, enhancing its practicality for NISQ devices. However, this advantage comes with costs: each iteration requires a canonical transformation of the Hamiltonian, which leads to an expanded Hamiltonian size. In practice, this requires high computational resources on classical computers, coupled to additional measurements on quantum devices, and gradient calculations at each step. Furthermore, the iQCC method does not fully guarantee the ordering of operators[43], which can increase the risk of encountering local minima and contribute to the ‘barren plateaus’ problem [44]. Nevertheless, recent advances have been made in exploring efficient techniques for QCC methods, such as in [45, 46].
@@ I.1 Disentangled unitary coupled cluster (DUCC): Lie Algebraic structure and symmetric excitations
where $\hat{P_{l}}$ is the so-called Pauli string operator, formed by the tensor product of Pauli matrices (with $l\in\left\{1,\ldots,M\right\}$), and $\vec{\theta}=\{\vec{\theta_{l}}\}_{l=1}^{M}$ are the variational parameters to be optimized. These operators do not necessarily commute, which can lead to a disordered wavefunction and significantly impact the results[51]. To address this ordering issue, we review the Lie algebraic properties of a specific set of operators below, ensuring they satisfy the closure relations. Moreover, since the main goal of the present work is to develop a compact ansatz suitable for NISQ devices, it is ideal to keep the set of operators as small as possible. Ideally, it should grow linearly or polynomially with the number of qubits, $n$. In practice, this can be achieved by applying symmetries to the multi-qubit space operators, which facilitates the optimization of the generators set. These symmetries include the particle number $\hat{N_{e}}$, spin symmetries ($\hat{S}^{2}=0$, $\hat{S_{z}}$ = 0), and symmetry point group (or spatial symmetry $\hat{R}$). We refer to these operators as “starters” since they act directly on top of the Hartree-Fock state, guaranteeing a symmetric initial wavefunction. This is, in turn, expected to accelerate convergence, as demonstrated in previous studies [37, 52]. Assume now we have a set of starters $\mathcal{S}_{j}=\{\hat{P}_{l}\}_{l=1}^{M_{s}}$ of length $M_{s}$, with $\hat{P}_{l}$ being symmetry-preserving operators. Then, it would be possible to build a sub-algebra from this set, denoted as $\mathcal{S}$, involving $M\leq 4^{n-1}$, satisfied by closure relations such that for any pair
@@ I.2 The NI-DUCC-VQE protocol
Figure 1: Schematic of NI-DUCC-VQE algorithm. (1) The necessary conditions for constructing a robust symmetry-preserving minimal complete pool (MCP) of size $22n-$ 2. The initial qubit excitations (starters) are strong and symmetric, effectively screened from fermionic excitations. A brief example demonstrates the symmetry operations on the qubit excitations in BeH2 molecule. (2) Once the MCP is generated, the NI-DUCC ansatz is constructed from layers $k$. To the left, an example of mapping the exponential of a Pauli operator into a circuit is shown, by using staircase method [7, 10, 20]. (3) Numerical results tested on BeH2 (14 qubits), using NI-DUCC-VQE, showing the comparison between robust and non-robust MCP excitations, in terms of chemical accuracy. (4) A list of applications that can benefit from NI-DUCC-VQE and its ansatz.
@@
    Refining the selection of double excitations through symmetries enforcement. To obtain the qubit Pauli string operators needed for constructing the NI-DUCC real wavefunction (see Eq.1), we create an odd Pauli string, $\hat{P}$, from each of the pre-selected fermionic operator $(i,j,k,l)$, which involves an odd number of $Y$ Pauli matrices. We then filter out the Pauli strings that violate the symmetry rules when acting on the Hartree-Fock state, such as $\hat{N_{e}}$, $\hat{S}^{2}$, $\hat{S}_{z}$ as well as the Abelian point group operator $\hat{R}$ (the reader should refer to the detailed examples of such operations presented in section 4 of the supplementary materials.[53]). The use of this technique ensures that the Pauli operators satisfy the condition $\langle\psi_{HF}|[\hat{H},\hat{P}_{l}]|\psi_{HF}\rangle\neq 0$. This means that the operators $\hat{P}_{l}$ are selected as initial operators because their commutator with the Hamiltonian is significantly non-zero. We denote these operators $\hat{P}_{l}$ as strongly symmetric qubit excitations.
@@ IV.2 Challenges in MCP generation as the number of qubits increases
Number of qubits Number of electrons Minimal Pool (MCP) Number of starters Number of FullSet Number of FullGroup
8 4 14 320 992 2048
10 5 18 $2.82\times 10^{3}$ $1.63\times 10^{4}$ $3.28\times 10^{4}$
12 6 22 $2.61\times 10^{4}$ $2.62\times 10^{5}$ $5.24\times 10^{5}$
14 7 26 $1.88\times 10^{5}$ $4.19\times 10^{6}$ $8.39\times 10^{6}$
16 8 30 $1.38\times 10^{6}$ $6.71\times 10^{7}$ $1.34\times 10^{8}$
18 9 34 $3.43\times 10^{7}$ $1.07\times 10^{9}$ $2.15\times 10^{9}$
20 10 38 $1.01\times 10^{9}$ $1.72\times 10^{10}$ $3.44\times 10^{10}$
22 11 42 $2.56\times 10^{10}$ $2.74\times 10^{11}$ $5.50\times 10^{11}$
@@ V Conclusion
We have introduced the NI-DUCC-VQE algorithm, designed to construct a disentangled unitary coupled-cluster ansatz, based on a specific set of qubit excitation generators. This method scales linearly with the number of qubits, and leverages strong symmetry orbitals along with Lie algebraic properties. On a test set of several molecules, we have demonstrated that our method systematically improves accuracy, achieving full convergence, and reaching the Full Configuration Interaction (FCI) reference energy. This is particularly significant at larger internuclear distances, where increased electron correlation leads to multiple electron configurations contributing comparably to the wavefunction. We have highlighted that this aspect is frequently ignored in the current state-of-the-art quantum computing fixed ansatz for quantum chemistry. This is true even when using higher-order fermionic excitations in the unitary coupled cluster (UCC) method, such as up to triple excitations (UCCSDT) or even up to quadruple excitations (UsCCSDTQ). Such methods fail to capture all the correlations, especially when molecular bond lengths deviate from equilibrium. NI-DUCC not only demonstrates an high level of accuracy but also proves to be a hardware-efficient approach, making it well-suited for implementation on NISQ devices. The CNOT count for the NI-DUCC ansatz scales linearly with its depth $k$, as $\mathcal{O}(knp)$, due to the $22n-$ 2 generators of MCP that scale linearly, which results in a relatively shallow circuit depth. When comparing NI-DUCC-VQE to ADAPT-VQE, particularly the promising QEB-ADAPT-VQE algorithm, which targets a shallower circuit depth, we found that similar CNOT counts can be achieved. NI-DUCC-VQE has also shown that with a small number of layers, it struggles to accurately find the exact ground state. However, with a sufficient number of layers, the algorithm consistently, and rapidly, converges to the FCI solution. We also observed that when the number of layers is sufficiently large, the energy decreases exponentially with each optimization step. This demonstrates that if the NI-DUCC ansatz incorporates more layers, the optimization hypersurface becomes more favorable. This effectively eliminates the false local minima, as demonstrated with the tested molecules in the article, thereby achieving FCI solution with fewer iterations. This finding paves the way for future discussions on using the NI-DUCC method to address the barren plateaus problem, where a cost function minimization reveals flat areas in the cost landscape of a variational quantum algorithm.
@@
Besides that, our NI-DUCC (Eq. 3) is structurally similar to the periodic structure ansatz (see for example Eq. (2) in [67]), since both rely on layer growth. Periodic structure ansatz have been extensively used to discuss the overparameterization phenomenon in quantum neural networks (QNN) [67]. Overparameterizing a neural network can enhance QNN performance, reduce training and generalization errors [68, 69], and even result in provable convergence [70, 71]. This similarity opens the doors to extend NI-DUCC into periodic structure ansatz problems, by integrating them with QNN frameworks in quantum chemistry. Furthermore, because the NI-DUCC wavefunction relies on Lie algebraic structures, which results in a linear scaling of CNOT gates, it could be beneficial to apply it to other variational quantum algorithms. The utility of Lie-algebraic structures has been demonstrated in efficient classical simulations across various variational quantum computing algorithms, such as the longitudinal-transverse field Ising model and the quantum approximate optimization algorithm (see details in Section III of ref. [72]). In fact, the dynamic lie (DL) algebraic theory [67, 72] has been employed there, to generate sub-algebras of dimension $\mathcal{O}(\text{poly}(n))$, where $n$ denotes the number of qubits [72]. It was found that understanding the theoretical connections between the DL sub-algebras and the gradients computations, as demonstrated in [73], has the potential to enhance the trainability of quantum circuits, and exemplify their efficient implementation in numerous paradigmatic tasks in variational quantum computing. In the near future, exploring the integration of the Lie algebraic properties used in our article with dynamic Lie sub-algebra holds promise for advancing the capabilities of Lie algebraic structures in efficiently simulating quantum chemistry problems.

### v2

@@ Abstract
Due to their non-iterative nature, fixed unitary coupled cluster (UCC) ansätze are attractive for performing quantum chemistry variational quantum eigensolver (VQE) computations as they avoid pre-circuit measurements on a quantum computer. However, achieving chemical accuracy for strongly correlated systems with UCC requires further inclusion of higher-order fermionic excitations beyond triples increasing circuit depth. We introduce $k$-NI-DUCC, a fixed and non-iterative disentangled unitary coupled cluster compact ansatz, based on specific ``$k^{\prime\prime}$ sets of “qubit” excitations, eliminating the needs for fermionic-type excitations. These elements scale linearly ($\mathcal{O}(n)$) by leveraging Lie algebraic structures, with $n$ being the number of qubits. The key excitations are screened through specific selection criteria, including the enforcement of all symmetries, to ensure the construction of a robust set of generators. NI-DUCC employs ``$k^{\prime\prime}$ products of the exponential of $\mathcal{O}(n)$- anti-Hermitian Pauli operators, where each single Pauli string has a length $p$. This results in a fewer two-qubit CNOT gates circuit scaling, $\mathcal{O}(knp)$, suitable for hardware implementations. Tested on LiH, H6 and BeH2, NI-DUCC-VQE achieves both chemical accuracy and rapid convergence even for molecules deviating significantly from equilibrium. It is hardware-efficient, reaching the exact full configuration interaction energy solution at specific layers, while reducing significantly the VQE optimization steps. If NI-DUCC-VQE effectively addresses the gradient measurement bottleneck of ADAPT-VQE-like iterative algorithms, the classical computational cost of constructing the $\mathcal{O}(n)$ set of excitations increases exponentially with the number of qubits. We provide a first implementation for constructing the generators’ set, able to handle up to 20 qubits and discuss the efficiency perspectives.
@@ Introduction
To address these issues, constructing an ansatz tailored to system-specific criteria offers a promising solution for achieving shallower circuits. The popular ADAPT-VQE iterative algorithm is a notable example of this strategy [32]. While capable of extracting electronic correlation energy using a small subset of operators, the operator selection process corresponds itself to a high computational overhead associated to gradients measurements. Despite subsequent efforts to minimize these costs, ADAPT-VQE versions [33, 34, 35, 36, 37, 38, 39, 40] continue to face gradient measurement challenges, especially as molecular size increases, while maintaining the necessary chemical accuracy. Similar to ADAPT-VQE methods, the Qubit coupled cluster (QCC) approach [41, 42] and its variants, including iterative Qubit coupled cluster (iQCC) [43], offer promising strategies for designing a compact wavefunction. In iQCC, generators are chosen from a specific set of operators, including all possible Pauli strings, to ensure energy reduction. These generators are sampled using first-order derivative calculations, resulting in a shallow circuit depth. Unlike ADAPT-VQE, the iQCC method avoids the need for full re-optimization of all parameters at each iteration, enhancing its practicality for NISQ devices. However, this advantage comes with costs: each iteration requires a canonical transformation of the Hamiltonian, which leads to an expanded Hamiltonian size. In practice, this requires high computational resources on classical computers, coupled to additional measurements on quantum devices, and gradient calculations at each step. Furthermore, the iQCC method does not fully guarantee the ordering of operators[43], which can increase the risk of encountering local minima and contribute to the ‘barren plateaus’ problem [44]. Nevertheless, recent advances have been made in exploring efficient techniques for QCC methods, such as in [45, 46]. In addition to the qubit coupled cluster (QCC) and ADAPT-VQE approaches, recent advancements such as the tiled unitary product state (tUPS) [47] and subspace projective quantum eigensolver (SPQE) [48] methods have significantly improved computational accuracy and the efficiency of quantum computing resources.
@@ I.1 Disentangled unitary coupled cluster (DUCC): Lie Algebraic structure and symmetric excitations
where $\hat{P_{l}}$ is the so-called Pauli string operator, formed by the tensor product of Pauli matrices (with $l\in\left\{1,\ldots,M\right\}$), and $\vec{\theta}=\{\vec{\theta_{l}}\}_{l=1}^{M}$ are the variational parameters to be optimized. These operators do not necessarily commute, which can lead to a disordered wavefunction and significantly impact the results[53]. To address this ordering issue, we review the Lie algebraic properties of a specific set of operators below, ensuring they satisfy the closure relations. Moreover, since the main goal of the present work is to develop a compact ansatz suitable for NISQ devices, it is ideal to keep the set of operators as small as possible. Ideally, it should grow linearly with the number of qubits, $n$. In practice, this can be achieved by applying symmetries to the multi-qubit space operators, which facilitates the optimization of the generators set. These symmetries include the particle number $\hat{N_{e}}$, spin symmetries ($\hat{S}^{2}=0$, $\hat{S_{z}}$ = 0), and symmetry point group (or spatial symmetry $\hat{R}$). We refer to these operators as “starters” since they act directly on top of the Hartree-Fock state, guaranteeing a symmetric initial wavefunction. This is, in turn, expected to accelerate convergence, as demonstrated in previous studies [37, 54]. Assume now we have a set of starters $\mathcal{S}_{j}=\{\hat{P}_{l}\}_{l=1}^{M_{s}}$ of length $M_{s}$, with $\hat{P}_{l}$ being symmetry-preserving operators. Then, it would be possible to build a sub-algebra from this set, denoted as $\mathcal{S}$, involving $M\leq 4^{n-1}$, satisfied by closure relations such that for any pair
@@ I.2 The NI-DUCC-VQE protocol
Figure 1: Schematic of NI-DUCC-VQE algorithm. (1) The necessary conditions for constructing a robust symmetry-preserving minimal complete pool (MCP) of size $22n-$ 2. The initial qubit excitations (starters) are strong and symmetric, effectively screened from fermionic excitations. (2) Once the MCP is generated, the NI-DUCC ansatz is constructed from layers $k$. To the left, an example of mapping the exponential of a Pauli operator into a circuit is shown, by using staircase method [7, 10, 20]. (3) Numerical results tested on BeH2 (14 qubits), using NI-DUCC-VQE, showing the comparison between robust and non-robust MCP excitations, in terms of chemical accuracy. (4) A list of applications that can benefit from NI-DUCC-VQE and its ansatz.
@@
    Refining the selection of double excitations through symmetries enforcement. To obtain the qubit Pauli string operators needed for constructing the NI-DUCC real wavefunction (see Eq.1), we create an odd Pauli string, $\hat{P}$, from each of the pre-selected fermionic operator $(i,j,k,l)$, which involves an odd number of $Y$ Pauli matrices. We then filter out the Pauli strings that violate the symmetry rules when acting on the Hartree-Fock state, such as $\hat{N_{e}}$, $\hat{S}^{2}$, $\hat{S}_{z}$ as well as the Abelian point group operator $\hat{R}$ (the reader should refer to the detailed examples of such operations presented in section 4 of the supplementary materials.[55]). The use of this technique ensures that the Pauli operators satisfy the condition $\langle\psi_{HF}|[\hat{H},\hat{P}_{l}]|\psi_{HF}\rangle\neq 0$. This means that the operators $\hat{P}_{l}$ are selected as initial operators because their commutator with the Hamiltonian is significantly non-zero. These excitations are of double-type in nature. Since the molecular Hamiltonian contains only single and double excitations, any excitations that create more than double excitations on top of the Hartree-Fock state would result in zero commutators. We denote the operators $\hat{P}_{l}$ as strongly symmetric qubit excitations.
@@ IV.2 Challenges in MCP generation as the number of qubits increases
(qubits,electrons) Minimal Pool (MCP) Number of starters Number of FullSet Number of FullGroup
(8,4) 14 320 992 2048
(10,5) 18 $2.82\times 10^{3}$ $1.63\times 10^{4}$ $3.28\times 10^{4}$
(12,6) 22 $2.61\times 10^{4}$ $2.62\times 10^{5}$ $5.24\times 10^{5}$
(14,7) 26 $1.88\times 10^{5}$ $4.19\times 10^{6}$ $8.39\times 10^{6}$
(16,8) 30 $1.38\times 10^{6}$ $6.71\times 10^{7}$ $1.34\times 10^{8}$
(18,9) 34 $3.43\times 10^{7}$ $1.07\times 10^{9}$ $2.15\times 10^{9}$
(20,10) 38 $1.01\times 10^{9}$ $1.72\times 10^{10}$ $3.44\times 10^{10}$
(22,11) 42 $2.56\times 10^{10}$ $2.74\times 10^{11}$ $5.50\times 10^{11}$
@@ V Conclusion
We have introduced the NI-DUCC-VQE algorithm, designed to construct a disentangled unitary coupled-cluster ansatz, based on a specific set of qubit excitation generators. This method scales linearly with the number of qubits, and leverages strong symmetry orbitals along with Lie algebraic properties. On a test set of several molecules, we have demonstrated that our method systematically improves accuracy, achieving full convergence, and reaching the Full Configuration Interaction (FCI) reference energy. This is particularly significant at larger internuclear distances, where increased electron correlation leads to multiple electron configurations contributing comparably to the wavefunction. We have highlighted that this aspect is frequently ignored in the current state-of-the-art quantum computing fixed ansatz for quantum chemistry. This is true even when using higher-order fermionic excitations in the unitary coupled cluster (UCC) method, such as up to triple excitations (UCCSDT) or even up to quadruple excitations (UsCCSDTQ). Such methods fail to capture all the correlations, especially when molecular bond lengths deviate from equilibrium. NI-DUCC not only demonstrates an high level of accuracy but also proves to be a hardware-efficient approach, making it well-suited for implementation on NISQ devices. The CNOT count for the NI-DUCC ansatz scales as $\mathcal{O}(knp)$ with its depth $k$, due to the $22n-$ 2 generators of MCP that scale linearly, which results in a relatively shallow circuit depth. When comparing NI-DUCC-VQE to ADAPT-VQE, particularly the promising QEB-ADAPT-VQE algorithm, which targets a shallower circuit depth, we found that similar CNOT counts can be achieved. NI-DUCC-VQE has also shown that with a small number of layers, it struggles to accurately find the exact ground state. However, with a sufficient number of layers, the algorithm consistently, and rapidly, converges to the FCI solution. We also observed that when the number of layers is sufficiently large, the energy decreases exponentially with each optimization step. This demonstrates that if the NI-DUCC ansatz incorporates more layers, the optimization hypersurface becomes more favorable. This effectively eliminates the false local minima, as demonstrated with the tested molecules in the article, thereby achieving FCI solution with fewer iterations. This finding paves the way for future discussions on using the NI-DUCC method to address the barren plateaus problem, where a cost function minimization reveals flat areas in the cost landscape of a variational quantum algorithm.
@@
Besides that, our NI-DUCC (Eq. 3) is structurally similar to the periodic structure ansatz (see for example Eq. (2) in [69]), since both rely on layer growth. Periodic structure ansatz have been extensively used to discuss the overparameterization phenomenon in quantum neural networks (QNN) [69]. Overparameterizing a neural network can enhance QNN performance, reduce training and generalization errors [70, 71], and even result in provable convergence [72, 73]. This similarity opens the doors to extend NI-DUCC into periodic structure ansatz problems, by integrating them with QNN frameworks in quantum chemistry. Furthermore, because the NI-DUCC wavefunction relies on Lie algebraic structures, which results in a linear scaling of CNOT gates, it could be beneficial to apply it to other variational quantum algorithms. The utility of Lie-algebraic structures has been demonstrated in efficient classical simulations across various variational quantum computing algorithms, such as the longitudinal-transverse field Ising model and the quantum approximate optimization algorithm (QAOA) (see details in Section III of ref. [74]). In fact, the dynamic lie (DL) algebraic theory [69, 74] has been employed there, to generate sub-algebras of dimension $\mathcal{O}(\text{poly}(n))$, where $n$ denotes the number of qubits [74]. It was found that understanding the theoretical connections between the DL sub-algebras and the gradients computations, as demonstrated in [75], has the potential to enhance the trainability of quantum circuits, and exemplify their efficient implementation in numerous paradigmatic tasks in variational quantum computing. In the near future, exploring the integration of the Lie algebraic properties used in our article with dynamic Lie sub-algebra holds promise for advancing the capabilities of Lie algebraic structures in efficiently simulating quantum chemistry problems.